16x^2-72x-40=0 can someone help me solve this?
Answer by checkley79(3341)

16x^2-72x-40=0
8(2x^2-9x-5)=0
8(2x+1)(x-5)=0
2x=-1
x=-1/2 ans.
x-5=0
x=5 ans.